# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v2

@@
Abstract
Accurately forecasting multiple future events within a given time horizon is crucial for finance, retail, social networks, and healthcare applications. Event timing and labels are typically modeled using Marked Temporal Point Processes (MTPP), with evaluations often focused on next-event prediction quality. While some studies have extended evaluations to a fixed number of future events, we demonstrate that this approach leads to inaccuracies in handling false positives and false negatives. To address these issues, we propose a novel evaluation method inspired by object detection techniques from computer vision. Specifically, we introduce Temporal mean Average Precision (T-mAP), a temporal variant of mAP, which overcomes the limitations of existing long-horizon evaluation metrics. Our extensive experiments demonstrate that models with strong next-event prediction accuracy can yield poor long-horizon forecasts and vice versa, indicating that specialized methods are needed for each task. To support further research, we release HoTPP [^1] https://github.com/ivan-chai/hotpp-benchmark, the first benchmark designed explicitly for evaluating long-horizon MTPP predictions. HoTPP includes large-scale datasets with up to 43 million events and provides optimized procedures for both autoregressive and parallel inference, paving the way for future advancements in the field.
1 Introduction
@@
The world is full of events. Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information. In the era of AI, it is crucial to develop methods capable of handling these complex data streams. We refer to this type of data as Event Sequences (ESs). Event sequences differ fundamentally from other data types. Unlike tabular data (Wang & Sun, 2022), ESs include timestamps and possess an inherent order. In contrast to time series data (Lim & Zohren, 2021), ESs are characterized by irregular time intervals and structured descriptions of each event. These differences necessitate the development of specialized models and evaluation practices.
Sequence modeling is the primary task in Event Sequences (ESs) domain. In its simplest form, each event is defined by its type and occurrence time, a framework commonly known as Marked Temporal Point Processes (MTPP) (Rizoiu et al., 2017). Some studies extend MTPP to incorporate additional data fields and model complex dependencies between them (McDermott et al., 2024). However, the majority of MTPP approaches, as well as their evaluation pipelines, primarily focus on predicting the next event based on historical data.
@@
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks. For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions. However, the challenges of autoregressive prediction in the context of MTPP have not been thoroughly explored. Another difficulty lies in evaluation. Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su & Hua, 2017). Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths. Still, the limitations of this metric have not been previously considered.
In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies. Our key contributions are as follows:
  1. 1.
    We demonstrate that widely used evaluation methods for MTPPs often overlook critical aspects of model performance. We demonstrate that simple rule-based baselines can sometimes outperform popular deep learning methods when evaluated using OTD.
  2. 2.
    We introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision. T-mAP evaluates variable length sequences within a specified time horizon, as illustrated in Figure 1. Unlike previous approaches, T-mAP accurately accounts for false positives and false negatives while being invariant to linear calibration. Additionally, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
  3. 3.
    Using our established methodology, we demonstrate that high next-event prediction accuracy does not necessarily translate into high-quality long-horizon forecasts; in many cases, our experiments show the opposite. This highlights the necessity of developing specialized models for the long-horizon prediction task.
  4. 4.
    We release HoTPP, a new open-source benchmark designed to facilitate long-horizon event sequence prediction research. HoTPP brings together datasets and methods from various domains, including financial transactions, social networks, healthcare, and recommender systems, greatly expanding the diversity and scale of data compared to prior benchmarks. Additionally, we offer an efficient inference algorithm necessary for large-scale evaluations.
@@
Figure 1: Temporal mAP (T-mAP) evaluation pipeline. Unlike previous methods, T-mAP evaluates sequences of variable lengths within the prediction horizon. It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
2 Related Work
@@
Marked Temporal Point Processes. A Marked Temporal Point Process (MTPP) is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels (Rizoiu et al., 2017). Common approaches to MTPP modeling focus on predicting the next event. A basic solution involves independently predicting the event time and type. A more advanced approach splits the original sequence into subsequences, one for each event type, and independently models each subsequence’s timing. Depending on the time-step distribution, the process is called Poisson or Hawkes.
@@
Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures. Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei & Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024). In this work, we evaluate MTPP models with both discrete and continuous time architectures. Unlike previous benchmarks, we also assess simple rule-based predictors and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021) and Next-K models from time series analysis (Lim & Zohren, 2021). For more details on MTPP modeling, refer to Appendix A.
@@
MTPP Evaluation. Previous benchmarks have primarily focused on medical data or traditional MTPP datasets. EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev & van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability. Early MTPP benchmarks, such as Tick (Bacry et al., 2017) and PyHawkes [^2] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks. While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and Next-K approaches. Furthermore, PoPPy does not evaluate long-horizon predictions at all. EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
@@
Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su & Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided. This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW). In our study, we address the limitations of OTD, a variant of the Wasserstein Distance, and propose a new metric called T-mAP. Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
3 Limitations of the Next-Event and OTD Metrics
@@
Figure 2: Comparison of OTD and T-mAP metrics. In example 1.a prefix evaluation of three events using OTD results in an incorrect, false negative. T-mAP addresses this issue by comparing all events within the time horizon, as shown in example 1.b. Additionally, OTD evaluates only the label with the maximum probability. From the OTD perspective, cases 2.a and 2.b have the same quality. In contrast, T-mAP evaluates the entire distribution, yielding a low score of 0.33 in case 2.a and a high score of 1 in case 2.b.
@@
Recent studies have advanced the measurement of horizon prediction quality by employing Optimal Transport Distance (OTD) (Mei et al., 2019). This metric is analogous to the edit distance between the predicted event sequence and the ground truth. Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$. These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$. The prediction is connected to the ground truth event if their types are equal. Denote $M(S_{p},S_{gt})$ a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$. OTD finds the minimum cost among all possible matchings:
$\mathrm{OTD}(S_{p},S_{gt})=\min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$ (1)
@@
OTD is computed between fixed-size prefixes, but this approach limits metric flexibility in assessing models with imprecise time step predictions, as illustrated in Fig. 2.1. If the model predicts events too frequently, the first $K$ predicted events will correspond to the early part of the ground truth sequence, resulting in false negatives. This outcome is inaccurate because allowing the model to predict more events would alter the evaluation. Conversely, if the time step is too large, the first $K$ predicted events will extend far beyond the horizon of the first $K$ target events, leading to a significant number of false positives. This is undesirable as it prevents the inclusion of additional ground truth events in the evaluation. Therefore, evaluating a dynamic number of events is essential to properly align with the ground truth’s time horizon.
@@
MTPP evaluation metrics can be categorized into two groups: those considering event indices and those assessing time and label prediction quality independently of actual positions. Next-event metrics rely on ordering, even when events share identical timestamps, making the order meaningless. While Optimal Transport Distance (OTD) itself is invariant to order, the extraction of sequences for comparison is influenced by event indices. For example, when OTD compares prefixes of length $K$, the algorithm must place $K$ correctly predicted events at the beginning of a sequence to minimize the OTD score. As a result, OTD, like next-event metrics, can depend on event ordering even when the order cannot be uniquely determined. This dependency is non-trivial and difficult to measure or control accurately.
@@
Another limitation of the OTD metric is its inability to evaluate the full predicted distribution of labels, as shown in Figure 2.2. OTD considers only the labels with the highest probability, ignoring the complete distribution. This makes it dependent on model calibration and limits the ability to assess performance across a broader range of event types, such as long-tail predictions. However, models typically predict probabilities for all classes, allowing for a more comprehensive assessment. Therefore, we aim to develop an evaluation metric that accurately captures performance across common and rare classes.
4 Temporal mAP: Prediction as Detection
@@
In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration. T-mAP is inspired by object detection methods from computer vision (Everingham et al., 2010), as illustrated in Fig. 1. The goal of object detection is to localize objects within an image and identify their types. In event sequences, we tackle a similar problem but consider the time dimension instead of horizontal and vertical axes. Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration. Therefore, we replace the intersection-over-union (IoU) similarity between bounding boxes with the absolute time difference.
@@ 4.1 Definition
4.2 Computation
Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP. This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker & Volgenant, 1988). The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
Figure 3: T-mAP computation pipeline.
@@ 4.2 Computation
$\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$ (3)

### v3

@@
Abstract.
Forecasting multiple future events within a given time horizon is essential for applications in finance, retail, social networks, and healthcare. Marked Temporal Point Processes (MTPP) provide a principled framework to model both the timing and labels of events. However, most existing research focuses on predicting only the next event, leaving long-horizon forecasting largely underexplored. To address this gap, we introduce HoTPP, the first benchmark specifically designed to rigorously evaluate long-horizon predictions. We identify shortcomings in widely used evaluation metrics, propose a theoretically grounded T-mAP metric, present strong statistical baselines, and offer efficient implementations of popular models. Our empirical results demonstrate that modern MTPP approaches often underperform simple statistical baselines. Furthermore, we analyze the diversity of predicted sequences and find that most methods exhibit mode collapse. Finally, we analyze the impact of autoregression and intensity-based losses on prediction quality, and outline promising directions for future research. The HoTPP source code, hyperparameters, and full evaluation results are available on GitHub [^1] https://github.com/ivan-chai/hotpp-benchmark.
copyright: acmlicensedjournalyear: 2025
1. Introduction
@@
Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information. In the era of AI, it is crucial to develop methods capable of handling these complex data streams. We refer to this type of data as Event Sequences (ESs).
Event sequences differ fundamentally from other data types. Unlike tabular data (Wang and Sun, 2022), ESs include timestamps and possess an inherent order. In contrast to time series data (Lim and Zohren, 2021), ESs are characterized by irregular time intervals and additional data fields, such as event types. These differences necessitate the development of specialized models and evaluation practices.
Irregularly spaced timestamps are commonly modeled via Temporal Point Processes (TPP) (Rizoiu et al., 2017). Assigning categorical labels to these events leads to the Marked Temporal Point Process (MTPP) formalism. Since most event-sequence data can be viewed as MTPPs augmented with additional fields (McDermott et al., 2024), accurately modeling event times and labels remains a central challenge in the broader event-stream (ES) domain.
@@
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events presents unique challenges that differ from traditional next-event prediction tasks (Xue et al., 2022; Zhou et al., 2025). For instance, autoregressive event sequence prediction involves applying the model to its own potentially erroneous predictions. However, the challenges of autoregressive prediction in the context of MTPP have not been thoroughly explored. Another difficulty lies in evaluation. Methods like Dynamic Time Warping (DTW) are typically unsuitable for event sequences due to strict ordering constraints (Su and Hua, 2017). Some works have applied Optimal Transport Distance (OTD), a variant of the Wasserstein distance, to compare sequences of predefined lengths (Mei et al., 2019; Xue et al., 2022, 2023). However, the limitations of this metric have not been previously considered.
@@
Figure 1. T-mAP evaluation pipeline. Unlike previous approaches, T-mAP evaluates sequences of variable lengths within the prediction horizon. It further improves performance assessment by analyzing label distributions rather than relying on fixed predictions.
In this work, we provide the first in-depth analysis of models and metrics for long-horizon event forecasting, establishing a rigorous evaluation framework and a baseline for MTPP studies. Our key contributions are summarized as follows:
  1. (1)
    We introduce HoTPP, the first open-source benchmark specifically designed for long-horizon event forecasting. By considering datasets and methods spanning finance, social networks, healthcare, and recommender systems, HoTPP significantly broadens the diversity and scale of data compared to previous benchmarks (Xue et al., 2023; McDermott et al., 2024; Xu, 2018). Moreover, we provide optimized training and inference algorithms to support large-scale studies.
  2. (2)
    We demonstrate that widely used evaluation metrics for MTPPs overlook critical aspects of model performance. In particular, they are sensitive to calibration and incorrectly count false positives and false negatives. To address these issues, we introduce Temporal mean Average Precision (T-mAP), a novel evaluation metric inspired by best practices in computer vision (Lin et al., 2014). Furthermore, we address a theoretical gap in computer vision by proving the correctness of the T-mAP computation algorithm.
  3. (3)
    Guided by our methodology, we evaluate the forecasting performance of a diverse set of approaches. Our findings indicate that methods excelling at next-event prediction often underperform in long-horizon scenarios, and vice versa. Moreover, in 6 out of 10 long-horizon tasks, simple statistical baselines either outperform or match modern MTPP models, highlighting the need for more effective solutions.
  4. (4)
    Finally, we quantify the entropy of predicted labels and demonstrate that most modern approaches tend to predict only the most frequent labels, suffering from the problem known as mode collapse. This issue severely restricts practical applicability and motivates further research into more accurate prediction strategies.
2. Related Work
@@
Marked Temporal Point Processes. A Marked Temporal Point Process (MTPP) is a stochastic process consisting of a sequence of pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ represent event times and $l_{i}\in\{1,\dots,L\}$ denote event type labels (Rizoiu et al., 2017). Common approaches to MTPP modeling focus on predicting the next event. A basic solution involves independently predicting the event time and type. A more advanced approach splits the original sequence into subsequences, one for each event type, and independently models each subsequence’s timing. Depending on the form of time-step distribution, the process is called Poisson or Hawkes.
@@
Over the last decade, the focus has shifted toward increasing model flexibility by applying neural architectures. Several works have employed classical RNNs (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019) and transformers (Zuo et al., 2020; Zhang et al., 2020), while others have proposed architectures with continuous time (Mei and Eisner, 2017; Rubanova et al., 2019; Kuleshov et al., 2024). In this work, we evaluate MTPP models with both discrete and continuous time architectures. Unlike previous benchmarks, we also assess simple rule-based baselines and popular methods from related fields, including GPT-like prediction models for event sequences (McDermott et al., 2024; Padhi et al., 2021). For more details on MTPP modeling, refer to Appendix A.
@@
MTPP Evaluation. Previous benchmarks have primarily focused on medical data or traditional MTPP datasets. EventStream-GPT (McDermott et al., 2024) and TemporAI (Saveliev and van der Schaar, 2023) consider only medical data and do not implement methods from the MTPP field, despite their applicability. Early MTPP benchmarks, such as Tick (Bacry et al., 2017) and PyHawkes [^2] https://github.com/slinderman/pyhawkes, implement classical machine learning approaches but exclude modern neural networks. While PoPPy (Xu, 2018) and EasyTPP (Xue et al., 2023) include neural methods, they do not consider rule-based and horizon prediction approaches, such as HYPRO (Xue et al., 2022) and Diffusion (Zhou et al., 2025). Furthermore, PoPPy does not evaluate long-horizon predictions at all. EasyTPP addresses this limitation to some extent by reporting the OTD metric, though it does not provide the corresponding evaluation code.
@@
Previous work also highlighted the limitations of Dynamic Time Warping (DTW) for event sequence evaluation (Su and Hua, 2017), showing that DTW’s strict ordering constraints are impractical and should be avoided. This work proposed an alternative, the Order-preserving Wasserstein Distance (OPW). In our study, we address the limitations of OTD, a variant of the Wasserstein Distance, and propose a new metric called T-mAP. Unlike OPW, T-mAP operates on timestamps rather than event indices and involves two hyperparameters, compared to three in OPW.
@@
Figure 2. Comparison of the OTD and T-mAP metrics. In the example 1.a, OTD’s prefix-based evaluation of three events erroneously produces a false negative, whereas T-mAP in example 1.b compares all events within the specified time horizon for a more accurate assessment. Moreover, OTD evaluates only the label with the highest probability, treating cases 2.a and 2.b as equivalent. By contrast, T-mAP considers the entire distribution, yielding a low score (0.33) for 2.a and a perfect score (1) for 2.b.
3. T-mAP Metric
In this section, we show that widely used evaluation metrics for long-horizon prediction often miscount false positives (FP) and false negatives (FN). We then introduce Temporal mean Average Precision (T-mAP), a novel, theoretically grounded metric that addresses these issues. In addition, T-mAP is invariant to linear calibration of model outputs, simplifying and stabilizing comparisons across different methods. Additional T-mAP experiments on sequences with highly irregular time steps are presented in Appendix H. Long-tail performance measurement is discussed in Appendix I.
3.1. Limitations of the Next-Event and OTD Metrics
@@
Recent studies have advanced the measurement of horizon prediction quality by employing OTD (Mei et al., 2019). This metric is analogous to the edit distance between the predicted event sequence and the ground truth. Suppose there is a predicted sequence $S_{p}=\{(t^{p}_{i},l^{p}_{i})\}_{i=1}^{n_{p}},0<t_{1}^{p}<t_{2}^{p}<\dots$, and a ground truth sequence $S_{gt}=\{(t^{gt}_{j},l^{gt}_{j})\}_{j=1}^{n_{gt}},0<t_{1}^{gt}<t_{2}^{gt}<\dots$. These two sequences form a bipartite graph $\mathcal{G}(S_{p},S_{gt})$. The prediction is connected to the ground truth event if their types are equal. Denote $M(S_{p},S_{gt})$ a set of all possible matches in the graph $\mathcal{G}(S_{p},S_{gt})$. OTD finds the minimum cost among all possible matchings:
(1) $\mathrm{OTD}(S_{p},S_{gt})=\\ \min\limits_{m\in M(S_{p},S_{gt})}\left[\sum\limits_{(i,j)\in m}|t^{p}_{i}-t^{gt}_{j}|+C_{del}U_{p}(m)+C_{ins}U_{gt}(m)\right],$
@@
OTD is computed over fixed-size prefixes, limiting its flexibility when models produce imprecise time-step predictions, as shown in Fig. 2.1. When a model predicts events too frequently, the first $K$ predicted events will correspond to the early part of the ground truth sequence, resulting in false negatives. This outcome is misleading because considering a longer predicted sequence would alter the evaluation. Conversely, if the predicted time step is too large, the first $K$ predicted events will extend far beyond the horizon of the first $K$ target events, leading to a significant number of incorrect false positives. Therefore, evaluating a dynamic number of events is essential to align with the ground truth’s time horizon properly.
Figure 3. T-mAP computation pipeline.
@@
Metrics for MTPP evaluation can be categorized into two groups: those relying on event indices and those that measure time and label prediction quality independently of event sequential positions. Next-event metrics depend on ordering, even if multiple events share the same timestamp, which makes the order meaningless. Although OTD is formally invariant to order, extracting sequences for comparison still depends on event indices. For instance, when OTD compares prefixes of length $K$, the first $K$ predicted events must align with the ground truth to minimize the OTD score. As a result, OTD, like next-event metrics, can be influenced by ordering even when it cannot be uniquely determined. This dependency is non-trivial and cannot be easily measured or controlled.
@@
Another limitation of the OTD metric is its inability to evaluate the full predicted distribution of labels, as shown in Fig. 2.2. OTD considers only the labels with the highest probability, ignoring the complete distribution. This makes it dependent on model calibration and limits the ability to assess performance across a broader range of event types, such as long-tail predictions. However, models typically predict probabilities for all classes, allowing for a more comprehensive assessment. Therefore, we aim to develop an evaluation metric that accurately captures performance across common and rare classes.
3.2. Temporal mAP: Definition and Computation
@@
In this section, we introduce a novel metric, Temporal mAP (T-mAP), which analyses all errors within a predefined time horizon, explicitly controls ordering, and is invariant to linear calibration. T-mAP is inspired by object detection methods from computer vision (Everingham et al., 2010), as illustrated in Fig. 1. Object detection aims to localize objects within an image and identify their types. In event sequences, we tackle a similar problem but consider the time dimension instead of horizontal and vertical axes. Unlike object detection, where objects have spatial size, each event in an MTTP is a point without a duration. Therefore, we replace the intersection-over-union (IoU) similarity between bounding boxes with the absolute time difference.
@@ 3.2.1. Definition
Figure 4. T-mAP dependency on the $\delta$ parameter. The dashed line indicates the selected value.
3.2.2. Computation
Finding the optimal matching independently for each threshold value $h$ is impractical; thus, we need a more efficient method to evaluate T-mAP. This section shows how to optimize T-mAP computation using an assignment problem solver, like the Jonker-Volgenant algorithm (Jonker and Volgenant, 1988). The resulting complexity of T-mAP computation is $\mathcal{O}(BN^{3})$, where $B$ is the number of evaluated sequences and $N=\max(n_{p},n_{gt})$ is the number of events within the horizon.
@@ 3.2.2. Computation
(3) $\forall h\forall m\in M(\mathcal{G})\exists m_{h}\in M(\mathcal{G}_{h}):m_{h}\subset m.$